OpenStudy (anonymous):

(6x^8-5x^3+8x^2+7)+(7x^7+3x^3-8x)(-5x^3+2x^2+8x-2)

OpenStudy (anonymous):

expand it first

OpenStudy (anonymous):

Add

OpenStudy (anonymous):

write in exponential format

OpenStudy (anonymous):

This is the format

OpenStudy (anonymous):

2x^3+8x^2+2x-4 = 0 — -3x^3+5x^2+3x = 0 2x^3+8x^2+2x-4 = 0 — -2x^3-8x^2-3x+1 = 0 -2x^3-8x^2+3x-1 = 0 — -2x^3+8x^2-3x+7 = 0 -2x^3+8x^2+3x-7 = 0 — 2x^3-8x^2-4x+3 = 0 2x^3-8x^2+4x-3 = 0 — 2x^3+8x^2-4x+9 = 0 2x^3+8x^2+4x-9 = 0 — -2x^3-8x^2-5x+6 = 0 -2x^3-8x^2+5x-6 = 0 — -2x^3+8x^2-6x+2 = 0 -2x^3+8x^2+6x-2 = 0 — 2x^3-8x^2-6x+8 = 0 2x^3-8x^2+6x-8 = 0 — 2x^3+8x^2-7x+4 = 0 2x^3+8x^2+7x-4 = 0 — -2x^3-8x^2-8x+1 = 0 -2x^3-8x^2+8x-1 = 0 — -2x^3+8x^2-8x+7 = 0 -2x^3+8x^2+8x-7 = 0 — 2x^3-8x^2-9x+3 = 0 2x^3-8x^2+9x-3 = 0 — 2x^3+8x^2-9x+9 = 0 2x^3+8x^2+9x-9 = 0 — 2x^3-8x^2-x+6 = 0 2x^3-8x^2+x-6 = 0 — 2x^3-8x+8 = 0 2x^3+8x-8 = 0 — -2x^3+9x^2-2x+3 = 0 -2x^3+9x^2+2x-3 = 0 — 2x^3-9x^2-2x+9 = 0 2x^3-9x^2+2x-9 = 0 — 2x^3+9x^2-3x+5 = 0 2x^3+9x^2+3x-5 = 0 — -2x^3-9x^2-4x+2 = 0 -2x^3-9x^2+4x-2 = 0 — -2x^3+9x^2-4x+8 = 0 -2x^3+9x^2+4x-8 = 0 — 2x^3-9x^2-5x+4 = 0 2x^3-9x^2+5x-4 = 0 — 2x^3-9x^2+6x = 0 2x^3+9x^2-6x = 0 — -2x^3-9x^2-6x+7 = 0 -2x^3-9x^2+6x-7 = 0 — -2x^3+9x^2-7x+3 = 0 -2x^3+9x^2+7x-3 = 0 — 2x^3-9x^2-7x+9 = 0 2x^3-9x^2+7x-9 = 0 — 2x^3+9x^2-8x+5 = 0 2x^3+9x^2+8x-5 = 0 — -2x^3-9x^2-9x+2 = 0 -2x^3-9x^2+9x-2 = 0 — -2x^3+9x^2-9x+8 = 0 -2x^3+9x^2+9x-8 = 0 — -2x^3-9x^2-x+5 = 0 -2x^3-9x^2+x-5 = 0 — 2x^3-9x+5 = 0 2x^3+9x-5 = 0 — -2x^3-x^2-2x+2 = 0 -2x^3-x^2+2x-2 = 0 — -2x^3+x^2-2x+8 = 0 -2x^3+x^2+2x-8 = 0 — 2x^3-x^2-3x+4 = 0 2x^3-x^2+3x-4 = 0 — 2x^3-x^2+4x = 0 2x^3+x^2-4x = 0 — -2x^3-x^2-4x+7 = 0 -2x^3-x^2+4x-7 = 0 — -2x^3+x^2-5x+3 = 0 -2x^3+x^2+5x-3 = 0 — 2x^3-x^2-5x+9 = 0 2x^3-x^2+5x-9 = 0 — 2x^3+x^2-6x+5 = 0 2x^3+x^2+6x-5 = 0 — -2x^3-x^2-7x+2 = 0 -2x^3-x^2+7x-2 = 0 — -2x^3+x^2-7x+8 = 0 -2x^3+x^2+7x-8 = 0 — 2x^3-x^2-8x+4 = 0 2x^3-x^2+8x-4 = 0 — 2x^3-x^2+9x = 0 2x^3+x^2-9x = 0 — -2x^3-x^2-9x+7 = 0 -2x^3-x^2+9x-7 = 0 — 2x^3+x^2-x+3 = 0 2x^3+x^2+x-3 = 0 — -2x^3-x+3 = 0 -2x^3+x-3 = 0 — -3x^2+2x = 0 +3x^2-2x = 0 — -3x^2-3x+3 = 0 -3x^2+3x-3 = 0 — +3x^2-4x+5 = 0 +3x^2+4x-5 = 0 — -3x^2-5x+8 = 0 -3x^2+5x-8 = 0 — -3x^2+7x = 0 +3x^2-7x = 0 — -3x^2-8x+3 = 0 -3x^2+8x-3 = 0 — +3x^2-9x+5 = 0 +3x^2+9x-5 = 0 — +3x^2-x+8 = 0 +3x^2+x-8 = 0 — 3x^3-2x^2+2x+2 = 0 3x^3+2x^2-2x-2 = 0 — 3x^3+2x^2+2x+8 = 0 -3x^3-2x^2-2x-9 = 0 — -3x^3-2x^2+3x+5 = 0 -3x^3+2x^2-3x-5 = 0 — -3x^3+2x^2+4x+1 = 0 3x^3-2x^2-4x-1 = 0 — 3x^3-2x^2+4x+7 = 0 3x^3+2x^2-4x-7 = 0 — 3x^3+2x^2+5x+3 = 0 -3x^3-2x^2-5x-4 = 0 — -3x^3+2x^2+6 = 0 3x^3-2x^2-6 = 0 — -3x^3+2x^2+6x+6 = 0 3x^3-2x^2-6x-6 = 0 — 3x^3-2x^2+7x+2 = 0 3x^3+2x^2-7x-2 = 0 — 3x^3+2x^2+7x+8 = 0 -3x^3-2x^2-7x-9 = 0 — -3x^3-2x^2+8x+5 = 0 -3x^3+2x^2-8x-5 = 0 — -3x^3+2x^2+9x+1 = 0 3x^3-2x^2-9x-1 = 0 — 3x^3-2x^2+9x+7 = 0 3x^3+2x^2-9x-7 = 0 — -3x^3+2x^2+x+4 = 0 3x

OpenStudy (anonymous):